OpenStudy (anonymous):

Find the slope of the line that passes through the points (2,-4), and (4,8). -6 1/6 6 -1/6

OpenStudy (jackellyn):

m= y2-y1/x2-x1 =(8-(-4))/(4-2) = 8+4/2 =12/2 m=?

OpenStudy (anonymous):

6?

OpenStudy (jackellyn):

Yep!

OpenStudy (anonymous):

thanks(:

OpenStudy (jackellyn):

np :)